OpenStudy (anonymous):

write a function for the situation and tell if it is contionous or discrete ? A movie store sells DVDs for 10$ each,what is the cost, C, of N DVDs ?

OpenStudy (anonymous):

C(X) = 10N?

OpenStudy (anonymous):

these are my options

OpenStudy (anonymous):

A. C=10 + N; discrete

OpenStudy (anonymous):

Sorry I meant C(N) = 10N, so for example: that if you buy 2 DVD's, then that is for $20.

OpenStudy (anonymous):

B. C=10 N; discrete C. c=10 + N ; contionous D. C=10 N ; Continoius

OpenStudy (anonymous):

which one would it be out of those 4 plz help its my math final for tomorrow

OpenStudy (anonymous):

I would go for C = 10N, because that is exactly what I said, but expressed in function terms: C(N) = 10N is the same as C = 10N

OpenStudy (anonymous):

D

OpenStudy (anonymous):

wait C or D ?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you so much